# SigFormer: Sparse Signal-Guided Transformer for Multi-Modal Human Action Segmentation

### v1

@@ 4.3. Implementation Details
Training Details. At the feature extraction stage, we set the feature dimension of IMU and human body key point to 300, the feature dimension of body bounding box points to 240, and the feature dimension of the state signals of scanner and printer to 60. We adopt Adam (Kingma and Ba, 2015) optimizer with the base learning rate of $1\times 10^{-4}$ with a $1\times 10^{-4}$ weight decay. The first 50 training epochs are set as a warm-up phase and the learning rate is halved gradually after every 50 epochs. We exclude the “Null” class from the evaluation when calculating the losses. The model is trained for 200 epochs for each split. We utilize the batch size of 32 during training.
@@ 4.4. Comparison with State-of-the-Art Methods
Table 1. Comparison with state-of-the-art methods on the OpenPack dataset. $I$ denotes the IMU data, $K$ denotes the human body key points, $B$ represent the human body bounding box points, $D$ denotes the state signals of IoT-enabled devices, and $+$ represents utilizing the information from multiple modalities. “(Late)” denotes the late fused version of the model with $I$ and $D$. “U0104”-“U0207” represent 6 independent test sets. Some methods have no results on “U0203”-“U0207” due to different dataset splits.
@@
Secondly, we compare SigFormer with other methods in the context of multi-modality inputs. SigFormer obtains an F1 score of 0.844 with the combination of “I+D”, demonstrating a 1.4% improvement over LTS-Net (Yoshimura et al., 2022b). Furthermore, SigFormer achieves an F1 score of 0.937 based on “I+K”, which is 2.0% higher than the method of Chai et al. (Chai et al., 2023). Notably, Uchiyama (Uchiyama, 2023) integrated models trained through five different methods. When employing the “I+K+B+D” modality combination, their first training method got an F1 score of 0.932. Adhering to identical training configurations, including modality types and data processing methodologies, SigFormer obtains the macro-average F1 score of 0.958, exhibiting a significant improvement of 2.6% compared to their method. In particular, SigFormer achieves the F1 score of over 0.9 on all six independent test sets, indicating its robustness for action segmentation. Under each of the same modality combinations, SigFormer exhibits superior performance. Even utilizing fewer modalities (“I+K”), SigFormer attains better results than Uchiyama (Uchiyama, 2023), which requires more modalities as input.
@@ 4.5. Ablation Study
To investigate the characteristics of each modality, we present the accuracy in action segmentation on each action class for each modality in Fig. 4. Among these modalities, We discover that human key points usually dominate the performance and achieve the highest accuracy on the majority of action classes, which is consistent with the results in Table 2. The reason may be that the human key points supply a wealth of spatial information, characterize the details of the relative position changes of human joints, and improve the discrimination of actions. The body bounding box points simply represent the person’s general location in space, and the spatial details are fewer than the human key points, resulting in poor performance. With motion information, IMU data is effective in distinguishing high-acceleration actions, such as “insert” and “put”. IMU data have the lowest performance (excluding state signals of devices), possibly because they only carry motion information and are unable to handle changes in spatial location and human-object interactions. On the other hand, state signals of IoT-enabled devices, i.e., the handheld scanner and label printer, are highly trustworthy sources of information for recognizing specific operations. As a result, these signals exhibit the highest accuracy in recognizing actions that have utilized the corresponding devices, such as “scanning label”. However, they are not effective in recognizing the majority of other actions.
@@
We try different combinations of multiple modalities and summarize the F1 scores in Table 2. Intuitively, the more modalities are involved, the better performance we obtain, since complementary information can be utilized to compensate missing features in the single modal data.
@@
Finally, when fusing all four modalities, we achieve the highest F1 score of 0.958. As shown in Fig. 4, the combination of all different modalities enables the complementarity of information between them, thereby obtaining an accurate representation of human actions.
In conclusion, the fusion of IMU data, human body key points data, and body bounding box data effectively captures the comprehensive action information, encompassing both motion stream and spatial stream. Furthermore, incorporating information from IoT-enabled devices further enhances the model’s performance, providing precise action cues.
@@
As shown in Fig. 5, we visualize the predictions of different fusion methods at the $softmax$ layer. The straightforward concatenation method exhibits category prediction errors over extended periods, indicating its limited utilization of complementary information from various modalities. While the predictions from BF blocks and standard cross-attention are generally accurate, instances of category confusion and instability still arise. Compared to other methods, our proposed sparse guided cross-modal module reduces the potential confusion between similar action categories. Of particular note is that SigFormer exhibits increased stability in action category predictions linked to sparse signals, specifically those categorized under class index 7 (“scanning label”). This observation underscores SigFormer’s efficacy in integrating multi-modal signals characterized by varying information densities to yield precise representations of human actions.
@@
$\mathcal{L}_{c_{final}}$ $\mathcal{L}_{b_{final}}$ $\mathcal{L}_{c_{inner}}$ $\mathcal{L}_{b_{inner}}$
@@
Inner Class Loss. As illustrated in Table 4, the performance of our method can be significantly improved by calculating the inner class loss for the three types of information-rich data (IMU data, human body key points, and body bounding box points). This demonstrates that the inner class loss introduces the enhanced awareness of action class for each modality, enabling effective learning of discriminative feature representations. However, computing inner class loss for state signals of IoT devices fails to enhance the performance. This is probable because the state signals are relatively temporally sparse and only associated with the specific actions, making it impossible to directly determine the whole action classes from this information alone. Moreover, as shown in Table 5, when inner class loss is included in the modality combinations, the model performs significantly better, which also illustrates the effectiveness of introducing class awareness for each modality.
Inner Boundary Loss. Similarly to the experiments of inner class loss, computing the inner boundary loss for the first three modalities produces the best performance for action segmentation (Table 4). This favorable outcome can be attributed to the fact that each feature individually learns temporal boundary information, thus providing accurate boundary awareness for subsequent action segmentation. In contrast, due to the sparse and indistinct temporal boundaries in the devices’ state signals information, calculating the inner boundary loss for it lowers the F1 score by 0.8%. Furthermore, as demonstrated in Table 5, the introduction of inner boundary loss is able to improve the F1 score by 1.1%, underscoring the efficacy of integrating boundary awareness within each dense signal.
@@ 5. CONCLUSION
In this paper, we introduce SigFormer, an efficient network for the action segmentation task utilizing both sparse and dense signals. The core component of SigFormer is the sparse guided cross-modal module, which adopts the sparse signals as the guidance to fuse signals with different information densities. It employs masked attention to constrain the attention operation on specific temporal regions. Moreover, to augment the network’s awareness of temporal boundaries, we develop the intermediate bottleneck module equipped with the inner loss functions, which extracts features in respect to both action category and temporal boundary. Furthermore, the mutual interactive branches are designed to explicitly model multi-stage interactions between action category and temporal boundary. In summary, SigFormer not only leads to improved fusion performance, but also mitigates the over-segmentation errors. Comprehensive experiments demonstrate the superior performance of SigFormer on the OpenPack dataset by achieving a remarkable F1 score.

### v2

@@ 4.3. Implementation Details
Training Details. At the feature extraction stage, we set the feature dimension of IMU and human body key points to 300, the feature dimension of body bounding box points to 240, and the feature dimension of the state signals of the scanner and printer to 60. We adopt the Adam (Kingma and Ba, 2015) optimizer with the base learning rate of $1\times 10^{-4}$ with a $1\times 10^{-4}$ weight decay. The first 50 training epochs are set as a warm-up phase and the learning rate is halved gradually after every 50 epochs. We exclude the “Null” class from the evaluation when calculating the losses. The model is trained for 200 epochs for each split. We utilize the batch size of 32 during training.
@@ 4.4. Comparison with State-of-the-Art Methods
Table 1. Comparison with state-of-the-art methods on the OpenPack dataset. $I$ denotes the IMU data, $K$ denotes the human body key points, $B$ represents the human body bounding box points, $D$ denotes the state signals of IoT-enabled devices, and $+$ represents utilizing the information from multiple modalities. “(Late)” denotes the late fused version of the model with $I$ and $D$. “U0104”-“U0207” represent 6 independent test sets. Some methods have no results on “U0203”-“U0207” due to different dataset splits.
@@
Secondly, we compare SigFormer with other methods in the context of multi-modality inputs. SigFormer obtains an F1 score of 0.844 with the combination of “I+D”, demonstrating a 1.4% improvement over LTS-Net (Yoshimura et al., 2022b). Furthermore, SigFormer achieves an F1 score of 0.937 based on “I+K”, which is 2.0% higher than the method of Chai et al. (Chai et al., 2023). Notably, Uchiyama (Uchiyama, 2023) integrated models trained through five different methods. When employing the “I+K+B+D” modality combination, their first training method got an F1 score of 0.932. Adhering to identical training configurations, including modality types and data processing methodologies, SigFormer obtains the macro-average F1 score of 0.958, exhibiting a significant improvement of 2.6% compared to their method.
Especially on the U0110, U0204, and U0207 test sets, compared to Uchiyama (Uchiyama, 2023), SigFormer demonstrates significant performance improvements, with an average enhancement of 5%. The average duration of actions in these test sets ranges from 7 to 8 seconds, whereas in other test sets, it spans from 10 to 13 seconds. Shorter action durations pose a more significant challenge for action segmentation. SigFormer incorporates action boundary information as constraints in both the feature extraction and final prediction phases. This approach enables accurate identification of temporal action boundaries, thereby enhancing the model’s ability to predict short action segments.
Moreover, SigFormer achieves the F1 score of over 0.9 on all six independent test sets, indicating its robustness for action segmentation. Under each of the same modality combinations, SigFormer exhibits superior performance. Even utilizing fewer modalities (“I+K”), SigFormer attains better results than Uchiyama (Uchiyama, 2023), which requires more modalities as input.
@@ 4.5. Ablation Study
To investigate the characteristics of each modality, we present the accuracy in action segmentation on each action class for each modality in Fig. 4. Among these modalities, We discover that human key points usually dominate the performance and achieve the highest accuracy on the majority of action classes, which is consistent with the results in Table 2. The reason may be that the human key points supply a wealth of spatial information, characterize the details of the relative position changes of human joints, and improve the discrimination of actions. The body bounding box points simply represent the person’s general location in space, and the spatial details are fewer than the human key points, resulting in poor performance. With motion information, IMU data is effective in distinguishing high-acceleration actions, such as “insert” and “put”. IMU data have the lowest performance (excluding state signals of devices), possibly because they only carry motion information and are unable to handle changes in spatial location and human-object interactions. On the other hand, state signals of IoT-enabled devices, i.e., the handheld scanner and label printer, are highly trustworthy sources of information for recognizing specific operations. As shown in Figure 4, these signals demonstrate the highest accuracy in recognizing actions that involve the corresponding devices, such as ”scanning label”. However, they are less effective in recognizing most other actions. Consequently, as indicated in Table 2, the F1 score for predicting all actions solely utilizing state signals of the handheld scanner and label printer is only 0.191.
@@
We try different combinations of multiple modalities and summarize the F1 scores in Table 2. Intuitively, the more modalities involved, the better performance we obtain, as complementary information can compensate for missing features in single-modal data.
@@
Finally, by fusing all four modalities, the model achieved the highest F1 score of 0.958. As shown in Fig. 4, the combination of all different modalities enables the complementarity of information between them, thereby obtaining an accurate representation of human actions.
In conclusion, the fusion of IMU data, human body key points data, and body bounding box data effectively captures the comprehensive action information, encompassing both the motion stream and spatial stream. Furthermore, incorporating information from IoT-enabled devices further enhances the model’s performance, providing precise action cues.
@@
As shown in Fig. 5, we visualize the predictions of different fusion methods at the $softmax$ layer. The straightforward concatenation method exhibits category prediction errors over extended periods, indicating its limited utilization of complementary information from various modalities. While the predictions from BF blocks and standard cross-attention are generally accurate, instances of category confusion and instability still arise. Compared to other methods, our proposed sparse guided cross-modal module reduces the potential confusion between similar action categories. Of particular note is that SigFormer exhibits increased stability in action category predictions linked to sparse signals, specifically those categorized under class index 6 (“scanning label”). This observation underscores SigFormer’s efficacy in integrating multi-modal signals characterized by varying information densities to yield precise representations of human actions.
@@
$\mathcal{L}_{c}^{final}$ $\mathcal{L}_{b}^{final}$ $\mathcal{L}_{c}^{inner}$ $\mathcal{L}_{b}^{inner}$
@@
Inner Class Loss. As illustrated in Table 4, the performance of our method can be significantly improved by calculating the inner class loss for the three types of information-rich data (IMU data, human body key points, and body bounding box points). This demonstrates that the inner class loss introduces the enhanced awareness of action class for each modality, enabling effective learning of discriminative feature representations. However, computing inner class loss for state signals of IoT devices fails to enhance the performance. This is probable because the state signals are relatively temporally sparse and only correspond to specific actions, making it impossible to directly determine the whole action classes from this information alone. Moreover, as shown in Table 5, when inner class loss is included in the modality combinations, the model performs significantly better, which also illustrates the effectiveness of introducing class awareness for each modality.
Inner Boundary Loss. Similarly to the experiments of inner class loss, computing the inner boundary loss for the first three modalities produces the best performance for action segmentation (Table 4). This favorable outcome can be attributed to the fact that each feature individually learns temporal boundary information, thus providing accurate boundary awareness for subsequent action segmentation. In contrast, due to the temporal sparsity of the device state signals, the action boundaries inferred from them are also sparse. Consequently, calculating the inner boundary loss based on these signals introduces a bias in the model’s perception of the correct action boundary, resulting in a performance degradation of 0.8%. Furthermore, as demonstrated in Table 5, the introduction of inner boundary loss can improve the F1 score by 1.1%, underscoring the efficacy of integrating boundary awareness within each dense signal.
@@ 5. CONCLUSION
In this paper, we introduce SigFormer, an efficient network for the action segmentation task utilizing both sparse and dense signals. The core component of SigFormer is the sparse guided cross-modal module, which adopts the sparse signals as the guidance to fuse signals with different information densities. It employs masked attention to constrain the attention operation on specific temporal regions. Moreover, to augment the network’s awareness of temporal boundaries, we develop the intermediate bottleneck module equipped with the inner loss functions, which extracts features related to both action category and temporal boundary. Furthermore, the mutual interactive branches are designed to explicitly model multi-stage interactions between action category and temporal boundary. In summary, SigFormer not only leads to improved fusion performance but also mitigates the over-segmentation errors. Comprehensive experiments demonstrate the superior performance of SigFormer on the OpenPack dataset by achieving a remarkable F1 score.
6. ACKNOWLEDGMENTS
This work was sponsored by Beijing Nova Program (NO. 20220484063), XDC02050200.